OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 72°?

OpenStudy (anonymous):

Recall that the exterior angles of a regular polygon add up to 360. So do the following: \[n = \frac{360}{m\angle Exterior \space Angle}\]Then plug it into this formula. \[Sum \space of \space interior \space angles = (n - 2)180\]

OpenStudy (campbell_st):

the exterior angles of a polygon sum to 360 the number of sides(angles) can be found by dividing 360 by the exterior angle 360/72 = 5 the polygon has 5 sides the to find the angle sum can be found using \[(n - 2)\times 180\]

OpenStudy (anonymous):

thank you to both @campbell_st and @Calcmathlete :D

OpenStudy (anonymous):

np :)